a sample of helium gas flows at a rate of 100cm3/s through a tube having a radius of 0.5cm. what is the speed of the He?
Answer #1

flow rate is given by:

Q = V*A

Q = flow rate of Helium gas = 100 cm^3/sec

V = Speed of He = ?

A = Cross-sectional Area of tube = pi*R^2

R = Radius of tube = 0.5 cm

So,

V = Q/A = Q/(pi*R^2)

Using given values:

V = (100 cm^3/sec)/(pi*0.5*0.5 cm^2)

V = 127.3 cm/sec = Speed of He

If you need final answer in m/sec, then

1 m = 100 cm, So

V = (127.3 cm/sec)*(1 m/100 cm)

V = 1.273 m/sec

Let me know if you've any query